OpenStudy (spareb665):

Solve the equation 2^3x+1=5^-x and round your answer to two digits after the decimal sign.

OpenStudy (zehanz):

Is it: \[2^{3x+1}=5^{-x}\]or\[2^{3x}+1=5^{-x}\]

OpenStudy (spareb665):



OpenStudy (zehanz):

Take the log on both sides

OpenStudy (zehanz):

Then use the rule: \[\log a^b=b \cdot \log a\]